OpenStudy (anonymous):

Need help understanding the concept: In group theory, the associativity rule says: m: G x G --> G, and i: G--> G where m: multiplication and i: inverse For every g, h, and k in G, m(m(g, h), k) =m (g, m(h, k) and for every g in G, m(g,i(g)) =e=m(i(g),g) Please explain me

OpenStudy (anonymous):

@ganeshie8 give me example, please

ganeshie8 (ganeshie8):

i have no clue sorry :( @zzr0ck3r

OpenStudy (anonymous):

Thanks but wonder why!! is it not group theory which you master ??

ganeshie8 (ganeshie8):

I have taken number theory before... i won't be useful here really sorry

OpenStudy (anonymous):

oh, number theory is different from group theory, hihihi... Thanks for reply.

OpenStudy (phi):

m: G x G --> G that says "m" takes two input, each from G, and produces an output which is in G m(m(g, h), k) =m (g, m(h, k) is notation that means the same thing as (g*h) * k = g * (h*k) (which is the associative property)

OpenStudy (anonymous):

oh, I go it, thank you very much. the same with the second one, Let me interpret it multiplication between g and its inverse = multiplication between g inverse and g, am I right?

OpenStudy (phi):

it's obviously more abstract than "times", because we don't (yet) have definition of "m", i.e. how it maps G x G to G

OpenStudy (phi):

Yes, but the second statement m(g,i(g)) =e=m(i(g),g) looks more like the commutative property. However, they are saying its the associative property, so there must be a way to interpret it that way.

OpenStudy (anonymous):

and e is identity element in the group. right?

OpenStudy (phi):

yes, it must be (by definition, a*i(a) = identity )

OpenStudy (anonymous):

oh, it makes sense now. Thank you so much.

OpenStudy (phi):

yw